OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. cos 107° cos 42° + sin 107° sin 42°

OpenStudy (anonymous):

sin 65° cos 65° cos 149° sin 149°

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

use the formula \(\Large \cos A\cos B + \sin A \sin B = \cos(A-B)\)

OpenStudy (anonymous):

I got 65

OpenStudy (anonymous):

so it would be cos65?

hartnn (hartnn):

107-42 = 65 so, yes! cos 65 is correct :)